# Translation from the Information Bottleneck Perspective: an Efficiency Analysis of Spatial Prepositions in Bitexts

### 20 Mar 2026 v1

@@ 1 Introduction
Languages partition semantic domains in remarkably diverse ways. For example, Levinson (2003) and Levinson et al. (2003) intensively document how spatial relations are named differently by speakers of typologically diverse languages. Yet beneath this diversity, systematic regularities emerge when naming spatial relations, but also within the semantic domains of colour and kinship inter alia (Kemp et al., 2018). One influential standpoint invokes the principle of efficient communication to explain these regularities (Gibson et al., 2019; Kemp et al., 2018; Zaslavsky et al., 2018). According to this view, when speakers of the same language balance the informational and cognitive load in their utterances, the mapping from meanings to words is optimised with respect to that trade-off between informativity and simplicity. Languages around the world achieve different realisations of this trade-off through diachronic evolution (Kirby et al., 2015).
The Information Bottleneck (IB) framework (Tishby et al., 2000) was put forward by Zaslavsky et al. (2018) to operationalise this trade-off for natural languages. It states that languages achieve near-optimal compression of meanings from an information-theoretic perspective, and has recently gained substantial empirical support. Usually, an IB analysis focuses on a specific semantic domain, and relies on naming data (e.g. natural descriptions of stimuli such as pictures or videos). The underlying assumption is that a stimulus evokes a mental representation, which is efficiently compressed into a description. The communicative cost of the description (complexity) is then compared to the relevant information it conveys about the stimuli (informativity), across natural and hypothetical languages. Prior IB applications relied on various stimuli like colour chips (Zaslavsky et al., 2018), static visual referents from images (Gualdoni et al., 2023), and dynamic kinematic events in videos of human locomotion (Langlois et al., 2025). However, to our knowledge, no IB analysis has considered linguistic stimuli.
@@
We tested this new methodology on a cornerstone of spatial language: spatial prepositions. These terms overtly mark for spatial relations between entities, and are highly prone to engage spatial representations (Landau and Jackendoff, 1993). We first extracted cross-linguistic variations of these spatial terms in four languages, namely detecting occurrences in source sentences from LIFAT et al. (2024), and aligning them with their target counterparts. This allowed to assess the complexity of spatial prepositions translations. Furthermore, to estimate the informativity of these translations, we also modelled the psychological space of spatial relations by leveraging pile-sorting results from a pilot study ($N=35$). We show that a low-dimensional space suffices to predict similarity judgements. Ultimately, we provide preliminary evidence for greater communicative efficiency of natural translations compared to random and counterfactual systems, in the semantic domain of spatial relations.
@@ 2 Theoretical framework
This section outlines how the IB theory has been applied to language. Tracing back to the Rate-Distortion theory (Cover, 1999), communication between a speaker and a listener is formalised as a joint search for an encoder and a decoder that enable lossy data compression. In this schematic communicative scenario, the speaker seeks a linguistic encoder $q$ that maps meanings to forms, which is minimally complex, and simultaneously minimise the distortion between the intended meaning $M$, and the listener’s reconstructed meaning $\hat{M}$. An operationalisation of this trade-off was first proposed by Tishby et al. (2000), and later applied to colour-naming (Zaslavsky et al., 2018), and other semantic domains (Mollica et al., 2021; Chen et al., 2023; Langlois et al., 2025).
[…]
@@ 3 General methodology
Within a specific semantic domain, an IB analysis relies on two types of data to compute the Complexity and Accuracy of an encoder $q$: naming data and human similarity judgements. Examples of naming data include colour names (when describing colour chips in Zaslavsky et al. (2018)), verbs of human gait (when describing a person moving in short videos, in Langlois et al. (2025)), object names (when describing real-world images in Gualdoni et al. (2023)), or spatial relations (when describing drawings of spatial scenes in Khetarpal et al. (2013)). In this work, we apply the IB framework to the translation task. In this section, we detail how naming data and similarity judgements can be recovered from translation data, and how Complexity and Accuracy measures are derived. Using a bitext, we treat each source sentence as a written stimulus, and specifically the term of interest $U$ as evoking the meaning $M$ for the human translator. The mental representation $M$ is then linguistically encoded as $W$, that is the translation of $U$. The schema $U\rightarrow M\rightarrow W$ summarises how translations are generated under these assumptions.
@@ 3.1 Naming data
Given bitexts such as those shown in Table 1, our goal is to align the terms of interest. For instance, we can align French spatial prepositions from the source text (e.g., $u_{1}=\textrm{devant}$, $u_{2}=\textrm{dans}$ and $u_{3}=\textrm{à l'est de}$) with their English counterparts in the target text ($w_{1}=\textrm{before}$, $w_{2}=\textrm{inside}$ and $w_{3}=\textrm{east of}$). This process yields alignment tables like the one presented in Table 2. From these alignments, we can derive the conditional probability distribution $p_{q}(w|m)$ where $m$ ranges over each distinct source term and $w$ ranges over the target lexicon. Assuming a uniform prior distribution $p(m)$ over all meanings in the source text  [^2] We use a uniform prior distribution as a first step., we compute the joint distribution: $p_{q}(w,m)=p_{q}(w|m)p(m)$. From this, we derive the Complexity $I_{q}(M;W)$ (see footnote 1 for details).
@@ 3.2 Similarity judgements
Collecting contextual similarity judgements of source terms grows quadratically in the size of the bitext. Being able to estimate the pairwise similarity between any pair of source terms is therefore crucial.
As part of our proposed methodology, we first select a subset of terms in their sentential context $\{u_{i}\mid 1\leq i\leq K\}$ of reasonable size $K$, and collect human pairwise similarity judgements $\{sim(u_{i},u_{j})\mid 1\leq i,j\leq K\}$. For ease of exposition, we denote these empirical similarities as $sim_{i,j}$.
Based on this annotated subset, we then train a model to predict the similarity $\widehat{sim}_{i,j}$ for any new unseen pair of terms. Following Zaslavsky et al. (2018) and Chen et al. (2023), we derive a subjective belief distribution on world states conditional on meanings $p(u|m)$, which is estimated as
@@
Here, $\gamma$ is a temperature parameter that controls the penalty for mistaking two distinct stimuli [^3] We set $\gamma$ so that $p(u|m)$ is defined by a softmax over similarity scores. Alternative methods for grounding the distribution $p(u|m)$ should be explored in future work.. Using the distributions $p(u|m)$, $p_{q}(w|m)$, and $p(m)$, we can now compute the joint distribution: $p_{q}(w,u)=p_{q}(w|m)p(m)p(u|m)$. This joint distribution allows us to derive $I_{q}(W;U)$.
@@ 4 Experiments
For the remainder of this work, we focus on the specific semantic domain of spatial relations. These relations are typically marked by prepositions as discussed extensively in the literature (for French linguistic aspects we refer to Vandeloise (1991); Borillo (1998); Le Pesant (2012), for a broad discussion on spatial cognition we refer to Landau and Jackendoff (1993)). We also consider ”absolute use” of prepositions (Le Pesant, 2012), which refers to instances where a preposition appears without its standard complement because the referent is understood from context (e.g., Je me gare devant / I park in front / Ich parke davor / Parkiram ispred). More importantly, we consider a preposition to have a spatial sense when its surrounding context triggers concrete spatial mental imagery, involving entities that can be physically touched or pointed at. The operational definition is theoretically grounded in the notions of cline of concreteness (Brysbaert et al., 2014) and cline of collocation (Pragglejaz, 2007).
Throughout our experiments [^4] Data and code is available at the following repository: https://github.com/antoineOYO/spatial-terms-translations, we utilised English, German, and Serbian translations of Jules Verne’s French adventure novel, Le tour du monde en 80 jours. The sentence-level alignment across these four texts was previously processed by LIFAT et al. (2024).
@@ 4.1 Naming data: detection and alignment of spatial prepositions
We first collected occurrences of phrases with potential spatial sense, guided by the inventory established by Le Pesant (2012). An initial disambiguation of each occurrence was carried out jointly by three annotators until consensus was reached, and 1,312 prepositions with a spatial sense were retained. Using the Large Language Model mistral-large-2512 (Mistral AI, 2025) in a few-shot setting, these French prepositions were subsequently aligned with their English, German and Serbian counterparts. The alignment table for Table 1 is illustrated in Table 2. The few-shot prompt for French-to-English alignment is available in Appendix A. To evaluate the quality of these automatic alignments, we sampled 100 preposition pairs for each target language. Errors of alignment were defined as the alignments of source prepositions to either inadequate target prepositions, or to any distinct other parts of speech, such as nouns or verbs. This evaluation yielded precision scores of 81% for French–English, 91% for French–German, and 94% for French–Serbian. Out of the 1,312 source prepositions, 580 were successfully aligned across all three target languages. The remaining instances were discarded, primarily because the translation did not rely on a preposition to encode the spatial relation in one or more target languages. The cross-linguistic discrepancy between automatic alignment is discussed in section 5.3.
Furthermore, we computed contextual embeddings for each source preposition, using xlm-roberta-large (Conneau et al., 2020). We averaged and stacked the last four hidden layers, yielding embeddings of size $d=4,096$. We then applied K-means clustering ($K=30$) to obtain a representative subset of 30 French sentences, each with a spatial preposition for downstream analysis. We found that $K=30$ reflected both the distribution of French spatial terms in the corpus, and the embedding space (see Appendix E for a dimension reduction). The clustering and dimension reduction were implemented using scikit-learn (Pedregosa et al., 2011).
@@ 4.2 Models of human similarity judgements
For this purpose, we designed a pile-sorting task based on the $K=30$ samples obtained from our clustering step. To assess the feasibility of these similarity judgements, we conducted an anonymised pilot study (see section 7) with a cohort of colleagues ($N=35$), all of whom were native French speakers. The reported patterns should be interpreted as preliminary evidence.
The 30 samples were printed on individual cardboard cards. The participants were orally instructed to group the samples into as many piles as they deemed necessary based on their perceived similarity; there was no time limit and participants could revise their groupings until satisfied. The full instructions (see Appendix B) were printed and available to each participants throughout the whole task. The distribution of number of piles produced by the participants is displayed in Appendix C.
Following Khetarpal et al. (2013), we define the pairwise similarity $sim_{i,j}$ as the proportion of all participants who sorted $u_{i}$ and $u_{j}$ into the same pile. A visualisation of the proportions is shown in Figure 3(a). From the list of pairwise similarities, a Multidimensional Scaling (MDS) [^5] Note that this MDS is for visualisation purpose only. was then computed using scikit-learn implementation (Figure 3(b)). Both visualisations suggest that the similarity structure of spatial relations is complex and graded rather than discrete or reducible to surface form.
@@
Figure 4: Attested (stars) and counterfactual (dots) translations of spatial prepositions. As the proportion of random permutations applied to the attested translations increases, counterfactual systems deviate further from the IB curve. Points are slightly jittered along the Complexity axis for visualisation.
@@ 4.3 Computation of encoders locations in the information plane
To benchmark the attested translations, we generated 3x10,000 counterfactual ”fake encoders” by randomly permuting 1%, 5% and 10% of the rows within the alignment tables. We also generated 100,000 uniform-random translations. We then computed the Accuracy and Complexity for both the attested and counterfactual translations using the implementation provided by Zaslavsky et al. (2018).
The optimal frontier was computed using the reversed deterministic annealing algorithm implementation from Chen et al. (2023), and was traced by iterating over 100 values of $\beta$ log-linearly spaced between 1 and $2^{20}$. Attested, counterfactual and optimal translations are displayed in Figure 4. We note that attested translations lie far from the steepest part of the theoretical bound, which is discussed in section 5.3. Since truly random translations yielded Accuracy scores much lower than the rest, they are not displayed in Figure 4 (average Accuracy was $2.6\pm 0.6\times 10^{-4}$ bits, average Complexity was $4.44\pm 0.16$ bits, a close-up of the information plane is displayed in Appendix D).
@@
The deviations of attested versus counterfactual translations is shown in Figure 5. Results indicate that as the degree of perturbation increases, the efficiency of translation decreases.
Figure 5: Deviation from optimality of attested translation, random and counterfactual ones. The black error bars stand for standard deviations.
@@ 5.1 Characterising cross-linguistic use of spatial prepositions
The automatic detection and alignment of spatial prepositions in bitexts has received limited attention until recent studies such as those by Viechnicki et al. (2024) and Beekhuizen (2025). For detection, we handled the bitext differently by choosing not to apply a syntactic filter to the source sentences. This choice is based on the observation that spatial terms appear in a variety of syntactic contexts, and range from single terms (”sur”, ”dans”…), to nominal locutions (”à bord du”, ”au milieu de”…). With respect to alignment, Viechnicki et al. (2024) rely on the GIZA++ word alignment tool (Och and Ney, 2003). In contrast, the RUIMTE model proposed by Beekhuizen (2025) avoids traditional word alignment, instead identifying spatial relation markers by calculating statistical associations. Direct performance comparisons between these methods are difficult, as these previous works evaluated highly morphologically complex languages. While our pipeline introduces a dependency on few-shot extraction tools, it generalises to any bitext without requiring morphologically complex language-specific alignment models.

### 05 Apr 2026 v2

@@ 1 Introduction
Languages partition semantic domains in remarkably diverse ways. For example, Levinson (2003) and Levinson et al. (2003) intensively document how spatial relations are named differently by speakers of typologically diverse languages. Yet beneath this diversity, systematic regularities emerge when naming spatial relations, but also within the semantic domains of colour and kinship inter alia (Kemp et al., 2018). One influential standpoint invokes the principle of efficient communication to explain these regularities (Gibson et al., 2019; Kemp et al., 2018; Zaslavsky et al., 2018). According to this view, when speakers of the same language balance the informational and cognitive load in their utterances, the mapping from meanings to words is optimised with respect to that trade-off between informativity and simplicity. Languages around the world achieve different realisations of this trade-off through evolution (Kirby et al., 2015).
The Information Bottleneck (IB) framework (Tishby et al., 2000) was put forward by Zaslavsky et al. (2018) to operationalise this trade-off for natural languages. It states that languages achieve near-optimal compression of meanings from an information-theoretic perspective, and has recently gained substantial empirical support. Usually, an IB analysis focuses on a specific semantic domain, and relies on naming data (e.g. natural descriptions of stimuli such as pictures or videos). The underlying assumption is that a stimulus evokes a mental representation, which is efficiently compressed into a description. The communicative cost of the description (complexity) is then compared to the relevant information it conveys about the stimuli (informativity), across natural and hypothetical languages. Prior IB applications relied on various stimuli like colour chips (Zaslavsky et al., 2018), static visual referents from images (Gualdoni et al., 2023), and kinematic events in videos of human locomotion (Langlois et al., 2025). However, to our knowledge, no IB analysis has considered linguistic stimuli.
@@
We tested this new methodology on a cornerstone of spatial language: spatial prepositions. These terms overtly mark for spatial relations between entities, and are highly prone to engage spatial representations (Landau and Jackendoff, 1993). We first extracted cross-linguistic variations of these spatial terms in four languages, namely detecting occurrences in source sentences from LIFAT et al. (2024), and aligning them with their target counterparts. This allowed us to assess the complexity of spatial prepositions translations. Furthermore, to estimate the informativity of these translations, we also modelled the psychological space of spatial relations by leveraging pile-sorting results from a pilot study ($N=35$). We show that a low-dimensional space suffices to predict similarity judgements. Ultimately, we provide preliminary evidence for greater communicative efficiency of natural translations compared to random and counterfactual systems, in the semantic domain of spatial relations.
@@ 2 Theoretical framework
This section outlines how the IB theory has been applied to language. Tracing back to the Rate-Distortion theory (Cover, 1999), communication between a speaker and a listener is formalised as a joint search for an encoder and a decoder that enable lossy data compression. In this schematic communicative scenario, the speaker seeks a linguistic encoder $q$ that maps meanings to forms, which is minimally complex, and simultaneously minimises the distortion between the intended meaning $M$, and the listener’s reconstructed meaning $\hat{M}$. An operationalisation of this trade-off was first proposed by Tishby et al. (2000), and later applied to colour-naming (Zaslavsky et al., 2018), and other semantic domains (Mollica et al., 2021; Chen et al., 2023; Langlois et al., 2025).
In this line of work, the Complexity of the stochastic encoder $q$ is quantified by the mutual information [^1] The mutual information is a general measure of how related two random variables are. It measures the reduction in uncertainty in one variable when observing the other: I(X;Y)=H(X)H(X|Y)=H(Y)H(Y|X)I(X;Y)=H(X)-H(X|Y)=H(Y)-H(Y|X), with $H(X)$ being the entropy of $X$. Note that the joint distribution $p(X,Y)$ is all we need to compute the mutual information: $I(X;Y)=\sum_{x,y}p(x,y)\log\frac{p(x,y)}{p(x)\cdot p(y)}$. It is a symmetric measure: $I(X;Y)=I(Y;X)$. $I_{q}(M;W)$ between two random variables: $M$ and WW [^2] We refer to random variables using upper-case letters, and realisations with lower-case letters.. The mental representations the speaker wishes to communicate are denoted by $M$, and $W$ represents the discrete linguistic forms they are mapped onto, according to the policy $p_{q}(w|m)$ of the encoder $q$. Conversely, the distortion between the speaker’s intended meaning $M$, and the listener’s reconstructed meaning $\hat{M}$, is measured using the expected Kullback-Leibler distance, $\mathbb{E}_{q}[D(M||\hat{M}$]. Minimising this distortion is mathematically equivalent to maximising the quantity $I_{q}(W;U)$, which is defined as the Accuracy of the encoder (Zaslavsky, 2019). In this formulation, $U$ represents the possible states of the world that the speaker perceives and describes, and about which the listener forms probabilistic inferences.
@@ 3 General methodology
Within a specific semantic domain, an IB analysis relies on two types of data to compute the Complexity and Accuracy of an encoder $q$: naming data and human similarity judgements. Examples of naming data include colour names (when describing colour chips in Zaslavsky et al. (2018)), verbs of human gait (when describing a person moving in short videos, in Langlois et al. (2025)), object names (when describing real-world images in Gualdoni et al. (2023)), or spatial relations (when describing drawings of spatial scenes in Khetarpal et al. (2013)). In this work, we apply the IB framework to the translation task. In this section, we detail how naming data and similarity judgements can be recovered from translation data, and how Complexity and Accuracy measures are derived.
[…]
@@ 3.1 Naming data
[…]
@@ 3.2 Similarity judgements
Collecting contextual similarity judgements of source terms grows quadratically in the size of the bitext. Being able to estimate the pairwise similarity between any pair of source terms is therefore crucial. As part of our proposed methodology, we first select a subset of terms in their sentential context $\{u_{i}\mid 1\leq i\leq K\}$ of reasonable size $K$, and collect human pairwise similarity judgements $\{sim(u_{i},u_{j})\mid 1\leq i,j\leq K\}$. For ease of exposition, we denote these empirical similarities as $sim_{i,j}$. Based on these pairwise similarity judgements, we then train a model to predict the similarity $\widehat{sim}_{i,j}$ for any new unseen pair of terms. Following Zaslavsky et al. (2018) and Chen et al. (2023), we derive a subjective belief distribution on world states conditional on meanings $p(u|m)$, which is estimated as
@@
Here, $\gamma$ is a temperature parameter that controls the penalty for mistaking two distinct stimuli [^4] We set $\gamma$ so that $p(u|m)$ is defined by a softmax over similarity scores. Alternative methods for grounding the distribution $p(u|m)$ should be explored in future work.. Note that when IB models a subjective belief as a distribution on world states $U$, it captures the ambiguity a term of interest can have in its context. Using the distributions $p(u|m)$, $p_{q}(w|m)$, and $p(m)$, we can now compute the joint distribution: $p_{q}(w,u)=p_{q}(w|m)p(m)p(u|m)$. From this, we derive the Accuracy $I_{q}(W;U)$.
@@ 4 Experiments
For the remainder of this work, we focus on the specific semantic domain of spatial relations. These relations are typically marked by prepositions as discussed extensively in the literature (for French linguistic aspects we refer to Vandeloise (1991); Borillo (1998); Le Pesant (2012), for a broad discussion on spatial cognition we refer to Landau and Jackendoff (1993)). We also consider “absolute use” of prepositions (Le Pesant, 2012), which refers to instances where a preposition appears without its standard complement because the referent is understood from context (e.g., Je me gare devant, I park in front, Ich parke davor and Parkiram ispred). More importantly, we consider a preposition to have a spatial sense when its surrounding context triggers a concrete spatial mental imagery, involving entities that can be physically touched or pointed at. This operational definition is theoretically grounded in the notions of cline of concreteness (Brysbaert et al., 2014) and cline of collocation (Pragglejaz, 2007).
Throughout our experiments [^5] Data and code are available at https://github.com/antoineOYO/spatial-terms-translations, we utilised English, German, and Serbian translations of Jules Verne’s French adventure novel, Le tour du monde en 80 jours. The sentence-level alignment across these four texts was previously processed by LIFAT et al. (2024).
@@ 4.1 Naming data: detection and alignment of spatial prepositions
We first collected occurrences of prepositions with potential spatial sense, namely searching for instances from the lexical inventory of spatial prepositions established by Le Pesant (2012). An initial disambiguation of each occurrence was then carried out jointly by three annotators until consensus was reached. The lexical and semantic disambiguation steps yielded 1,312 prepositions with a spatial sense. Using the Large Language Model mistral-large-2512 (Mistral AI, 2025) in a few-shot setting, these French prepositions were subsequently aligned with their English, German and Serbian counterparts. The alignment table for Table 1 is illustrated in Table 2. The few-shot prompt for French-to-English alignment is available in Appendix A. To evaluate the quality of these automatic alignments, we sampled 100 preposition pairs for each target language. Errors of alignment were defined as the alignments of source prepositions to either inadequate target prepositions, or to any distinct other parts of speech, such as nouns or verbs. This evaluation yielded precision scores of 81% for French–English, 91% for French–German, and 94% for French–Serbian. Out of the 1,312 source prepositions, 580 were successfully aligned across all three target languages. The remaining instances were discarded, primarily because the source preposition was not relexified with a preposition in at least one of the translations [^6] This is also known as zero-marking: consider the use of the preposition in Je suis à la maison versus I’m home. We refer to Beekhuizen (2025) for related analysis.. The cross-linguistic discrepancy between automatic alignments is discussed in section 5.3.
Furthermore, we computed contextual embeddings for each source preposition, using xlm-roberta-large (Conneau et al., 2020). We averaged and stacked the last four hidden layers, yielding embeddings of size $d=4,096$. We then applied $K$-means clustering ($K=30$) to obtain a representative subset of 30 French sentences, each with a spatial preposition, for downstream analysis. We found that $K=30$ reflected both the distribution of French spatial terms in the corpus, and the embedding space (see Appendix D for a dimension reduction). The clustering and dimension reduction were implemented using scikit-learn (Pedregosa et al., 2011).
@@ 4.2 Models of human similarity judgements
For this purpose, we designed a pile-sorting task based on the $K=30$ samples obtained from our clustering step. To assess the feasibility of these similarity judgements, we conducted an anonymised pilot study (see section 7) with a cohort of colleagues ($N=35$), all of whom were native French speakers. The reported patterns should be interpreted as preliminary evidence. The 30 samples were printed on individual cardboard cards. The participants were orally instructed to group the samples into as many piles as they deemed necessary based on their perceived similarity; there was no time limit and participants could revise their groupings until satisfied. The full instructions (see Appendix B) were printed and available to each participants throughout the whole task. The distribution of number of piles produced by the participants is displayed in Appendix C.
Following Khetarpal et al. (2013), we define the pairwise similarity $sim_{i,j}$ as the proportion of all participants who sorted $u_{i}$ and $u_{j}$ into the same pile. A visualisation of the proportions is shown in Figure 3(a). From the list of pairwise similarities, a Multidimensional Scaling (MDS) [^7] Note that this MDS is for visualisation purpose only. was then computed using scikit-learn implementation (Figure 3(b)). Both visualisations suggest that the similarity structure of spatial relations is complex and graded rather than discrete or reducible to surface form (discussed in section 5.2).
@@
(a) Information plane of attested, perturbed and random encoders of spatial prepositions. Dots are jittered slightly along the Complexity axis for visualisation purposes.
(b) Deviation from optimality for attested, perturbed and random encoders. Black error bars indicate standard deviations.
Figure 4: How much do attested and random systems deviate from optimality?
@@ 4.3 Computation of encoders locations in the information plane
To benchmark the attested translation encoders, we generated 3x10,000 counterfactual “perturbed encoders” by randomly permuting 1%, 5% and 10% of the rows within the alignment tables. We also generated 300,000 uniform-random translation encoders. We then computed the Accuracy and Complexity for attested, perturbed and fully random translation encoders using the implementation provided by Zaslavsky et al. (2018). The optimal frontier was computed using the reversed deterministic annealing algorithm implementation from Chen et al. (2023), and was traced by iterating over 100 values of $\beta$ log-linearly spaced between 1 and $2^{20}$. Attested, perturbed and random translations, along with the optimal frontier, are displayed in the information plane from Figure 4(a). We observe that attested translation encoders lie far from the steepest part of the theoretical bound, which is discussed in section 5.3.
@@
The deviations of translation encoders is shown in Figure 4(b). Results indicate that as the degree of perturbation increases, the efficiency of translation encoders decreases.
@@ 5.1 Characterising cross-linguistic use of spatial prepositions
The automatic detection and alignment of spatial prepositions in bitexts has received limited attention until recent studies such as those by Viechnicki et al. (2024) and Beekhuizen (2025). For detection, we handled the bitext differently by choosing not to apply a syntactic filter to the source sentences. This choice is based on the observation that spatial terms appear in a variety of syntactic contexts, and range from single terms (sur, dans), to nominal locutions (à bord du, au milieu de). With respect to alignment, Viechnicki et al. (2024) rely on the GIZA++ word alignment tool (Och and Ney, 2003). In contrast, the RUIMTE model proposed by Beekhuizen (2025) avoids traditional word alignment, instead identifying spatial relation markers by calculating statistical associations. Direct performance comparisons between these methods are difficult, as these previous works evaluated highly morphologically complex languages. While our pipeline introduces a dependency on few-shot extraction tools, it generalises to any bitext without requiring morphologically complex language-specific alignment models.